OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12 A. y = 11/9x + 4/3 slope: 4/3 y-intercept: 11/9 B. y = -11/9x + 4/3 slope: 11/9 y-intercept: 4/3 C. y = 11/9x - 4/3 slope: 11/9 y-intercept: 4/3 D. y = 11/9x - 4/3 slope: 11/9 y-intercept: -4/3

OpenStudy (ranga):

Slope-intercept form is: y = mx + b where m is the slope and b is the y-intercept. Rearrange -11x + 9y = -12 to look like y = mx + b. That is, isolate y on the left side and take everything else to the right. Simplify. Identify m and b.